OpenStudy (anonymous):

Help If f(x)=2x, then f^-1(x)= -2x ½x x - 2

OpenStudy (ness9630):

Switch x and y and solve for y y=2x ==> x=2y

OpenStudy (psymon):

We can say f(x) is the same as y. So basically we have y = 2x. To find the inverse, switch x and y and then resolve for y. So if we switch x and y we have x = 2y Now all we do is solve for y, meaning divide both sides by 2 to get y = x/2 or as you put, (1/2)x

OpenStudy (anonymous):

Thank you guys :)

OpenStudy (psymon):

Mhm ^_^

OpenStudy (ness9630):

Not surprised Psymon ;)

OpenStudy (austinl):

So, if you wanted to solve for \(f(x) = 2x+3\) \(f^{-1}(x)=?\)

OpenStudy (psymon):

If you wanted to do that, we would first switch x and y, meaning we would have: x = 2y + 3 Now as above, you would now solve for y. So subtracrting both sides by 3 and then dividing by 2, we get (x-3)/2 = y @austinL

OpenStudy (ness9630):

No need to teach Austin, Psymon

OpenStudy (psymon):

Why not? He asked, I have no reason not to answer.

OpenStudy (austinl):

It was directed at the asker to test their knowledge.

OpenStudy (ness9630):

I think he was asking Jess xD

OpenStudy (psymon):

Well, provided he only wanted an answer, I hope he doesnt mind that I answered instead xD

OpenStudy (austinl):

But yeah, I am slightly past this stage in mathematics :P

OpenStudy (psymon):

Fair enough xD

OpenStudy (ness9630):

Slightly? I think diff eq qualifies more than slightly

OpenStudy (psymon):

Haha, that, I think Diff Eqs is in another league from this xD

OpenStudy (austinl):

It really is a whole 'nother beast really.

OpenStudy (psymon):

Im in diff eq now, haha.